01line te3 Clem question C C O coo ngegenawcomilni akeAss Jnr Make cevelen Activi y doi ocatCT assignment-take5 akeAssigrmentsesso Locatc assignment-take Chapte 7 HW the Refer impurtant val fur this qu Desigu a buffer that has a pIT of 11.10 usirg one of week base conjugale acid systems shown below. Benf 4 But 2.1x10 10.62 a ton 6.7x10 S.17 be combined with bow may grams of ttc wcak basc, to produce 1.00 L of a buffer that How many grams of the chloride salt of the conjugate acid mur the weal, base? grantus chloride salt of cotjugale acid grans weak bases Submit Answer Retry Entire Group 6 more group attempts remaining Due Mar 24 at 50 PM nish Assignment Ask me 3rytning 1.00 M in Save and Di 6 PM 3/24/2011
Answer #1

pH = 11.10

pOH = 14-11.1 = 2.90

pKb nearest to this is CH3NH2 so:

pOH = pKB + log(BH+/B)

pKB = -log(4.2*10^-4) = 3.3767

2.90 = 3.3767 + log(CH3NH3Cl / CH3NH2)

CH3NH3Cl / CH3NH2 = 10^(2.90-3.3767 ) = 0.333656

and we know

CH3NH3Cl +CH3NH2 = 1

solve

CH3NH2 = 1-CH3NH3Cl

CH3NH3Cl / (1-CH3NH3Cl)= 0.333656

CH3NH3Cl = 1 - 0.333656*CH3NH3Cl

1.333656*CH3NH3Cl = 1

CH3NH3Cl = 1/1.333656 = 0.74981

so

CH3NH2 = 1-CH3NH3Cl

CH3NH2 = 1-0.74981 =0.25019

so

mass CH3NH2   = mol*MW = 0.25019*31.0571 = 7.7701 g

mass ofCH3NH3Cl = mol*MW = 0.74981*67.5180 = 50.625 g